OpenStudy (anonymous):

-5/4n=30 a- n= (-50) b- n= (-32) c- n= 32 d- n= 50

jigglypuff314 (jigglypuff314):

you want to get n all alone to find n from -5/4 n = 30 so multiply both sides by 4 *4 *4 so -5 n = 120 then divide both sides by negative 5 /(-5) /(-5) n = - 120/5 = ...